OpenStudy (winsteria):

MEDAL AND FAN A porch has the shape of a semicircle with a diameter of 7 ft. What is the approximate area of the porch, rounded to the nearest square foot? Use 3.14 to approximate pi star.

OpenStudy (winsteria):

@AlexandervonHumboldt2 @iGreen @StudyGurl14

OpenStudy (anonymous):

Firstly, do you know how to calculate area of a cirlce?

OpenStudy (anonymous):

a semicircle is just half of a circle.

OpenStudy (anonymous):

so just use. A = (pi*r^2)/2

OpenStudy (winsteria):

@arjund12 what's next?

OpenStudy (anonymous):

nothing, that is your answer.

OpenStudy (winsteria):

what's the steps to get my answer? @arjund12

OpenStudy (winsteria):

@iGreen

OpenStudy (anonymous):

diamete = 7 radius = 7/2 A = [(3.14)*(7/2)^2]/2 <- this is the area of the porch.

OpenStudy (anonymous):

A = 19.23

OpenStudy (anonymous):

ft^2

OpenStudy (winsteria):

19.2325?

OpenStudy (winsteria):

I have to round to the nearest whole number though?

OpenStudy (anonymous):

so 19 is the nearest whole number.

OpenStudy (winsteria):

Thank You

OpenStudy (anonymous):

No problem. :D